OpenStudy (anonymous):

Tom receives allowance every week. He spends three-fourths of his allowance on food. Last week, Tom spent nine dollars on food. What is Tom’s weekly allowance? Write the verbal model, define the variable, write the equation and find the solution. Show your work.

OpenStudy (anonymous):

Set up and equation. 3/4x = 9, where x represents his total allowance... Multiply each side by the reciprocal of the fraction to get rid of it. The reciprocal of 3/4 is 4/3. x = 9(4/3). x = 12.

OpenStudy (anonymous):

thank you :)